# 2Astex Pharmaceuticals, Cambridge CB4 0QA, UK

### v4

@@ 1.1 Cryo-EM Single Particle Analysis
Determining the structure of macromolecules is crucial to deciphering the intricacies of biological processes and the underlying mechanisms of diseases. With the advent of the resolution revolution, Cryogenic Electron Microscopy (Cryo-EM) has emerged as a leading technique for elucidating structures [35, 6]. This revolution, propelled by significant advances in electron direct detectors and image processing algorithms, has positioned Cryo-EM as a routine, often unrivaled, method for many complex samples [12]. Its advantages include, among others, the relative ease of sample preparation compared to other techniques (e.g., x-ray crystallography), the capability to analyze protein complexes previously considered out of reach, and the ability to recover different conformations, offering a dynamic view of molecules in action [36]. The pivotal role of Cryo-EM in structural biology was globally recognised in 2017 when the technique was awarded the Nobel Prize in Chemistry.
@@ 2.1 Benchmark Compilation and preprocessing
In our effort to build a comprehensive benchmark, our primary goal was to identify a diverse set of EMPIAR entries containing at least 200,000 particles, a number deemed sufficient for effective model training. Due to the limitations of EMPIAR’s search functionality and the inconsistencies in dataset annotations, we conducted a manual search for entries exceeding this particle count and containing standard Relion files ( .star and .mrcs). Subsequently, we verified the consistency and accuracy of the metadata by running relion_reconstruct [45] and visually assessing the resulting volumes. This step was crucial for eliminating a significant number of entries due to metadata issues that either crashed the reconstruction process or led to incorrect volumes. To ensure consistent estimation of particle poses, the data was then re-processed using Relion’s auto-refine [45]. Only entries for which the reconstructed volume exhibited resolution values close to those reported in the literature were selected for inclusion in the benchmark. Finally, for consistency, all images were downsampled to $1.5$/pixel, with different image dimensions in each entry as macromolecules vary in size (see Appendix A).
The results computed in this work were obtained with particles that were preprocessed on-the-fly. We performed per-image normalisation following the standard Cryo-EM procedure, which involves rescaling the intensity so that the background (noise) has a mean of 0 and a standard deviation of 1. We also corrected the contrast inversion caused by the defocus via phase flipping [39]. Finally, since the macromolecule typically represents only between 25% to 50% of the whole particle image, the images were cropped so that neighbouring particles are not included. It is important to note that our benchmark package allows users the flexibility to choose whether or not to apply any of these normalisation steps.
The data labels are represented as rotation matrices and then converted into grid indices by finding the closest rotation matrix in the $\mathrm{SO(3)_{grid}}$. For the cases in which the macromolecule exhibits point symmetry, the labels are expanded as $L_{i}={\{g_{j}R_{i}|g_{j}\in G}\}$ where $G$ is the set of rotation matrices given a point symmetry group (e.g., $C1$), and $R_{i}$ the ground-truth rotation matrix. As a result, the labels consist of a vector with $|G|$ non-zero values and $|\mathrm{SO(3)_{grid}}|-|G|$ zeros.
@@ 2.3 Evaluation metrics
The angular error ($\mathrm{angError}$) measures the geodesic distance between the predicted and ground-truth poses, typically expressed in degrees or radians. This distance can be directly calculated from the rotation matrix of the ground truth pose $\mathrm{trueR}_{i}$ and the predicted rotation matrix $\mathrm{predR}_{i}$ as
@@
which weights the $\mathrm{angError_{i}}$ by $\mathrm{conf_{i}}$, the confidence in the ground-truth pose, measured as Relion’s rlnMaxValueProbDistribution. This confidence estimation is a number between 0 and 1 that measures the probability of the particle having the reported ground truth orientation according to the Relion model. While $\mathrm{wMAnE}$ is still sensitive to ground truth and confidence estimation errors, due to its simplicity, we used it as the criterion for hyperparameter tuning.
The quality of volumes reconstructed from the predicted poses is assessed by comparing them with the ground-truth volumes generated from the original poses (see Appendix H). For this comparison, we employ the real space Pearson’s Correlation Coefficient (PCC) and the Fourier Shell Correlation (FSC) Resolution as metrics.
@@
In order to decouple the different quality levels of the different benchmark entries, we report the differences of the metrics with respect to the ground-truth levels, estimated from the half-maps of the ground truth,
@@
where $GT$ is the ground-truth map, $GT_{i}$ is the ground-truth map reconstructed with the $i$-th half of the data, and $V$ is the 3D volume reconstructed with the predicted poses (see Figure 1).
@@ 2.4 Training
where $Q(R_{c,i})$ is the predicted probability for the rotation matrix with grid index $c$ and $P(R_{c,i})$ is $1/|G|$ when any of the ground-truth matrices is $c$ and zero otherwise.
@@ 2.5 Evaluation protocol
Figure 1: Evaluation protocol inspired by the Cryo-EM Gold Standard. For each entry, the dataset is randomly split into two subsets (Half data 0 and Half data 1) that are processed independently. Then, each half of the data is used to train a different model that will be used to infer the poses of the other half of the data. From the inferred poses, two reconstructed volumes can be obtained, one for each half of the data. The two reconstructed volumes can be combined and compared to the ground-truth volume, which is obtained from the ground-truth poses. The grey box represents the automatic evaluation tool, that takes as input the predictions for both data half sets and internally computes the required reconstructed volumes to perform the comparisons.
Due to the uncertainty in the ground truth labels and the fact that what matters to Cryo-EM practitioners is the quality of the reconstructed volume, we devised an evaluation protocol inspired by the Cryo-EM gold standard [20], which is a per-entry 2-fold cross-validation strategy in which the poses of each half of the data are independently estimated and used to reconstruct two volumes (half-maps). For benchmarking supervised methods, it involves training an independent model for each half of the dataset to infer the poses of the other half of the dataset. After that, the final 3D volume is computed by reconstructing the two half-maps and averaging them. The final averaged map can then be compared with the ground-truth map obtained from the original poses (see Figure 1). It is important to note that the FSC resolution values derived from this comparison are analogous to map-to-model FSC resolution estimations and not equivalent to the gold standard half-to-half resolution.
Since Image2Sphere predicts only rotation matrices but not image shifts, when reconstructing the volumes, we employed the ground-truth translations. This could result in an overoptimistic estimation of performance, however, since the effect of the translations is tightly coupled with the accuracy of the angular estimation, this overestimation should be small. We leave for future work the full inference of both the rotations and translations. Finally, to avoid overfitting to the validation set, we performed hyperparameter tuning only one half of the data using $\mathrm{wMAnE}$ as a metric.
@@ 3.2 Performance of the baseline model on the benchmark
Table 1 summarises the results of the Image2Sphere [24] model on our benchmark, with per-entry results in Appendix D. While the $\mathrm{wMAnE}$ is $\sim$ 24°, for the best cases, the error is as small as 9°. The $\Delta\mathrm{PCC}$ for the worse cases is > 0.1, highlighting that, for some entries, the reconstructed volumes are far from the ground truth solution. For a few cases, the results are much better, with $\Delta\mathrm{PCC}$ < 0.03. In terms of prediction vs ground truth $\mathrm{FSCR_{0.5}}$, most maps are in the 8-6 Å range, with $\Delta\mathrm{FSCR_{0.5}}$ of 3.6 Å. However, the $\mathrm{FSCR_{0.143}}$ values between 4-5 Å, indicate better correlation at lower signal levels. This visually translates into a relatively well-resolved central part of the map that becomes blurrier away from the centre (see Figure 2 and Appendix E). For the top-performing cases, a simple and fast local refinement of the predicted poses is sufficient to obtain high-resolution reconstructions comparable to ground-truth volumes, at a computational cost threefold less than global refinement (Appendix F). Since the Image2Sphere model inference takes only minutes, far less than the hours needed for traditional refinement, further improvements could reduce computing times by at least one order of magnitude if local refinement is no longer needed (see Appendix G for running times).
Table 1: Image2Sphere results on CESPED. $\mathrm{MAnE}$ and $\mathrm{wMAnE}$ measure angular errors in predicted poses. $\Delta\mathrm{PCC}$ and $\Delta\mathrm{FSCR}$ measure the reduction in quality of the predicted volumes compared to the ground-truth. The mean and std of the metrics are computed over the seven benchmark entries.
@@
Figure 2: Ground-truth reconstruction for EMPIAR-10786 (grey) and reconstruction using the angles predicted with the Image2Sphere model (yellow).
@@ 3.3 Example of model generalisability across samples
Figure 3 showcases three reconstructed volumes: (1) EMPIAR-10278 using ground-truth poses (grey); (2) EMPIAR-10278 with poses predicted by the model trained on EMPIAR-10280 (yellow), illustrating the model’s generalisability; and (3) EMPIAR-10280 using poses inferred by the model trained on its own dataset, serving as a control for model performance. As expected, the EMPIAR-10278 map reconstructed with original poses shows superior quality compared to the others. Similarly, the EMPIAR-10280 map generated from the model trained on EMPIAR-10280 exhibits better quality than the EMPIAR-10278 map inferred using the EMPIAR-10280 model, reflecting the differences between the two datasets despite containing the same protein. Independently of these quality differences, the model’s capacity for generalization across datasets is evident through visual inspection of the EMPIAR-10278 inferred map (yellow), as the overall shape of the protein and several key secondary structure elements are clearly recognizable. This suggests that further improvements in the model could lead to the desired goal of training the model once and then inferring the poses of similar datasets at much faster speeds.
Figure 3: Generalisability use case. Left and centre, reconstructions for the EMPIAR-10278 dataset from the ground-truth poses (grey) and the predicted poses (yellow) using a model trained on the EMPIAR-10280 data. Right, reconstruction for the EMPIAR-10280 dataset using a model trained on the EMPIAR-10280 data used as control (cyan). The model used for the yellow and cyan volumes is the same, but the particles fed to the model come from different datasets of the same protein in different conditions. Although of lower quality than the cyan volume, the yellow volume demonstrates similarities to the ground-truth map (grey).
@@ 6 Acknowledgments
Appendix Appendix A Benchmark composition.
@@
Notice that rotation transformations require adjustments in the ground-truth labels.
@@
This section contains per-entry statistics for the Image2Sphere model predictions using the evaluation protocol proposed in the main text. These metrics compare the ground-truth map ($\mathrm{GT}$) against the map reconstructed from the predicted poses ($\mathrm{V}$), namely In particular, the $\mathrm{PCC}(GT,V)$ and $\mathrm{FSCR_{t}}(GT,V)$, where $t$ denotes the threshold 0.5 or 0.143, where reported. GT is obtained by employing relion_reconstruct on the ground truth poses (that were estimated with relion_refine --auto_refine). The reconstructed map V is generated with relion_reconstruct from the predicted poses.
@@
  • $\mathrm{MAnE}$: Mean Angular Error; $\mathrm{wMAnE}$: weighted Mean Angular Error; $\mathrm{PCC(V_{0},V_{1})}$: Reconstructed half-to-half Pearsons’s Correlation Coefficient; $\mathrm{PCC}(GT,V)$: Reconstructed to ground-truth Pearson’s Correlation Coefficient; $\mathrm{FSCR_{0.143}(V_{0},V_{1})}$: Reconstructed half-to-half FSC resolution at threshold 0.143, and $\mathrm{FSCR_{0.5}(V_{0},V_{1})}$ at threshold 0.5; $\mathrm{FSCR_{0.143}(GT,V)}$: Reconstructed to ground-truth resolution at threshold 0.143, and $\mathrm{FSCR_{0.5}}(GT,V)$ at threshold 0.5; $\mathrm{FSCR_{0.143}(GT_{0},GT_{1})}$: Ground-truth half-to-half FSC resolution at threshold 0.143, and $\mathrm{FSCR_{0.5}(GT_{0},GT_{1})}$ at threshold 0.5. $\mathrm{PCC}(GT_{0},GT_{1})$: Ground-truth half-to-half Pearsons’s Correlation Coefficient.
@@
This appendix shows the volumes reconstructed for some of the best performing examples of the Image2Sphere model on our benchmark. As it is shown in all the cases, the quality of the central region of the protein is quite close to the one of the ground-truth. However, the density for the regions that are at the edges of the macromolecule is much worse. This is in line to what could be expected if there were some degree of inaccuracy in the angular estimation, as the magnitude of the errors in the volume is proportional to both the angular error and the radius of the macromolecule.
Figure 1: Ground-truth reconstruction for EMPIAR-10280 (grey) and reconstruction using the angles predicted with the Image2Sphere model (yellow).
Figure 2: Ground-truth reconstruction for EMPIAR-10374 (grey) and reconstruction using the angles predicted with the Image2Sphere model (yellow).
@@
In this section, we illustrate the usefulness of our approach by showing the effect of classical local refinement on the Image2Sphere results for the benchmark entry 10374. In this case, the Image2Sphere model predicted poses with a $\mathrm{wMAnE}$ of 24.8° that lead to a reconstructed map with $\mathrm{FSCR_{0.143}}(V_{0},V_{1})$ of 3.7 Å. When the predicted poses are used as priors for a local refinement in Relion with --sigma_angle 2, the refined map achieved a $\mathrm{PCC}(GT,V)$ of 0.997 compared to the original 0.968, showing that the refined map is much more similar to the ground truth map. Indeed, as it can be appreciated in Supplementary Figure 3, after the local refinement, not only the quality of the core of the protein is comparable to the quality of the ground truth, but also the quality of the distant parts of the maps is much better, almost as good as in the ground truth. Equally important, since we are limiting the angular search to the neighbourhood around the predicted poses ($\pm$ 6°), the number of image comparisons carried out by Relion is much smaller, resulting in a three-fold speed-up in computational time, even when including the time required for pose inference using the Image2Sphere model.
Figure 3: Ground-truth reconstruction for EMPIAR-10374 (grey) and reconstruction using the angles predicted with the Image2Sphere model and locally refined using Relion with priors (yellow).
@@
Figure 5: Effect of particle misalignment in FSC resolution for entries 10647, 10786, and 11120. Left, FSC curves at different amounts of incorrectly aligned particles: 0% (blue), 10% (orange), 20% (green), and 30% (red) of the particles were assigned random angles. Right, FSC curves at different levels of induced angular inaccuracy. Each particle alignment was perturbed using uniform random noise of $\pm$0º (blue), 1º (orange), 3º (green), and 5º (red).

### v5

@@ 1.1 Cryo-EM Single Particle Analysis
Determining the structure of macromolecules is crucial to deciphering the intricacies of biological processes and the underlying mechanisms of diseases. With the advent of the resolution revolution, Cryogenic Electron Microscopy (Cryo-EM) has emerged as a leading technique for elucidating structures [36, 6]. This revolution, driven by significant advances in direct electron detectors and image processing algorithms, has made Cryo-EM a routine, often unrivaled, method for many complex samples [12]. Its advantages include, among others, the relative ease of sample preparation compared to other techniques (e.g., x-ray crystallography), the capability to analyze protein complexes previously considered out of reach, and the ability to recover different conformations, offering a dynamic view of molecules in action [37]. The pivotal role of Cryo-EM in structural biology was globally recognised in 2017 when the technique was awarded the Nobel Prize in Chemistry.
@@ 2.1 Benchmark Compilation and preprocessing
In our effort to build a comprehensive benchmark, our primary goal was to identify a diverse set of EMPIAR entries containing at least 200,000 particles, a number deemed sufficient for effective model training. Due to the limitations of EMPIAR’s search functionality and the inconsistencies in dataset annotations, we conducted a manual search for entries exceeding this particle count and containing standard Relion files ( .star and .mrcs). Subsequently, we verified the consistency and accuracy of the metadata by running relion_reconstruct [46] and visually assessing the resulting volumes. This step was crucial for eliminating a significant number of entries due to metadata issues that either crashed the reconstruction process or led to incorrect volumes. To ensure consistent estimation of particle poses, the data was re-processed using the Relion version 4 auto-refine program [46, 24] (see Appendix Appendix A for details). Only entries for which the reconstructed volume exhibited resolution values close to those reported in the literature were selected for inclusion in the benchmark. Finally, for consistency, all images were downsampled to $1.5$/pixel, with different image dimensions in each entry as macromolecules vary in size. See Appendix A for a list of the entries and their properties.
The images fed to the deep learning model were preprocessed on-the-fly. We performed per-image normalisation following the standard Cryo-EM procedure, which involves rescaling the intensity so that the background (noise) has a mean of 0 and a standard deviation of 1. We also corrected the contrast inversion caused by the defocus via phase flipping [40]. Finally, since the macromolecule typically represents only between 25% to 50% of the whole particle image, the images were cropped so that neighbouring particles are not included. It is important to note that our benchmark package allows users the flexibility to choose whether or not to apply any of these normalisation steps.
The data labels are represented as rotation matrices and then converted into grid indices by finding the closest rotation matrix in the $\mathrm{SO(3)_{grid}}$. For the cases in which the macromolecule exhibits point symmetry, the labels are expanded as $L_{i}={\{g_{j}R_{i}|g_{j}\in G}\}$ where $G$ is the set of rotation matrices given a point symmetry group (e.g., $C1$), and $R_{i}$ the ground truth rotation matrix. As a result, the labels consist of vectors with $|G|$ non-zero values and $|\mathrm{SO(3)_{grid}}|-|G|$ zeros.
@@ 2.3 Evaluation metrics
The angular error ($\mathrm{angError}$) measures the geodesic distance between the predicted and ground truth poses, typically expressed in degrees or radians. This distance can be directly calculated from the rotation matrix of the ground truth pose $\mathrm{trueR}_{i}$ and the predicted rotation matrix $\mathrm{predR}_{i}$ as
@@
which weights the $\mathrm{angError_{i}}$ by $\mathrm{conf_{i}}$, the confidence in the ground truth pose, measured as Relion’s rlnMaxValueProbDistribution. This confidence estimation is a number between 0 and 1 that measures the probability of the particle having the reported ground truth orientation according to the Relion model. While $\mathrm{wMAnE}$ is still sensitive to ground truth and confidence estimation errors, due to its simplicity, we used it as the criterion for hyperparameter tuning.
The quality of volumes reconstructed from the predicted poses is assessed by comparing them with the ground truth volumes generated from the original poses (see Appendix H). For this comparison, we employ the real space Pearson’s Correlation Coefficient (PCC) and the Fourier Shell Correlation (FSC) Resolution as metrics.
@@
In order to decouple the different quality levels of the different benchmark entries, we report the differences of the metrics with respect to the ground truth levels, estimated from the half-maps of the ground truth,
@@
where $GT$ is the ground truth map, $GT_{i}$ is the ground truth map reconstructed with the $i$-th half of the data, and $V$ is the 3D volume reconstructed with the predicted poses (see Figure 1).
@@ 2.4 Training
where $Q(R_{c,i})$ is the predicted probability for the rotation matrix with grid index $c$ and $P(R_{c,i})$ is $1/|G|$ when any of the ground truth matrices is $c$ and zero otherwise.
@@ 2.5 Evaluation protocol
Figure 1: Evaluation protocol inspired by the Cryo-EM Gold Standard. For each entry, the dataset is randomly split into two subsets (Half data 0 and Half data 1) that are processed independently. Then, each half of the data is used to train a different model that will be used to infer the poses of the other half of the data. From the inferred poses, two reconstructed volumes can be obtained, one for each half of the data. The two reconstructed volumes can be combined and compared to the ground truth volume, which is obtained from the ground truth poses. The grey box represents the automatic evaluation tool, that takes as input the predictions for both data half sets and internally computes the required reconstructed volumes to perform the comparisons.
Due to the uncertainty in the ground truth labels and the fact that what matters to Cryo-EM practitioners is the quality of the reconstructed volume, we devised an evaluation protocol inspired by the Cryo-EM gold standard [20], which is a per-entry 2-fold cross-validation strategy in which the poses of each half of the data are independently estimated and used to reconstruct two volumes (half-maps). For benchmarking supervised methods, it involves training an independent model for each half of the dataset to infer the poses of the other half of the dataset. After that, the final 3D volume is computed by reconstructing the two half-maps and averaging them. The final averaged map can then be compared with the ground truth map obtained from the original poses (see Figure 1). It is important to note that the FSC resolution values derived from this comparison are analogous to map-to-model FSC resolution estimations and not equivalent to the gold standard half-to-half resolution.
Since Image2Sphere predicts only rotation matrices but not image shifts, when reconstructing the volumes, we employed the ground truth translations. This could result in an overoptimistic estimation of performance, however, since the effect of the translations is tightly coupled with the accuracy of the angular estimation, this overestimation should be small. We leave for future work the full inference of both the rotations and translations. Finally, to avoid overfitting to the validation set, we performed hyperparameter tuning only one half of the data using $\mathrm{wMAnE}$ as a metric.
@@ 3.2 Performance of the baseline model on the benchmark
Table 1 summarises the results of the Image2Sphere [25] model on our benchmark, with per-entry results in Appendix D. While the $\mathrm{wMAnE}$ is $\sim$ 24°, for the best cases, the error is as small as 9°. The $\Delta\mathrm{PCC}$ for the worse cases is > 0.1, highlighting that, for some entries, the reconstructed volumes are far from the ground truth solution. For a few cases, the results are much better, with $\Delta\mathrm{PCC}$ < 0.03. In terms of prediction vs ground truth $\mathrm{FSCR_{0.5}}$, most maps are in the 8-6 Å range, with $\Delta\mathrm{FSCR_{0.5}}$ of 3.6 Å. However, the $\mathrm{FSCR_{0.143}}$ values between 4-5 Å, indicate better correlation at lower signal levels. This visually translates into a relatively well-resolved central part of the map that becomes blurrier away from the centre (see Figure 2 and Appendix E). For the top-performing cases, a simple and fast local refinement of the predicted poses is sufficient to obtain high-resolution reconstructions comparable to ground truth volumes, at a computational cost threefold less than global refinement (Appendix F). Since the Image2Sphere model inference takes only minutes, far less than the hours needed for traditional refinement, further improvements could reduce computing times by at least one order of magnitude if local refinement is no longer needed (see Appendix G for running times).
Table 1: Image2Sphere results on CESPED. $\mathrm{MAnE}$ and $\mathrm{wMAnE}$ measure angular errors in predicted poses. $\Delta\mathrm{PCC}$ and $\Delta\mathrm{FSCR}$ measure the reduction in quality of the predicted volumes compared to the ground truth. The mean and std of the metrics are computed over the seven benchmark entries.
@@
Figure 2: Ground truth reconstruction for EMPIAR-10786 (grey) and reconstruction using the angles predicted with the Image2Sphere model (yellow).
@@ 3.3 Example of model generalisability across samples
Figure 3 showcases three reconstructed volumes: (1) EMPIAR-10278 using ground truth poses (grey); (2) EMPIAR-10278 with poses predicted by the model trained on EMPIAR-10280 (yellow), illustrating the model’s generalisability; and (3) EMPIAR-10280 using poses inferred by the model trained on its own dataset, serving as a control for model performance. As expected, the EMPIAR-10278 map reconstructed with original poses shows superior quality compared to the others. Similarly, the EMPIAR-10280 map generated from the model trained on EMPIAR-10280 exhibits better quality than the EMPIAR-10278 map inferred using the EMPIAR-10280 model, reflecting the differences between the two datasets despite containing the same protein. Independently of these quality differences, the model’s capacity for generalization across datasets is evident through visual inspection of the EMPIAR-10278 inferred map (yellow), as the overall shape of the protein and several key secondary structure elements are clearly recognizable. This suggests that further improvements in the model could lead to the desired goal of training the model once and then inferring the poses of similar datasets at much faster speeds.
Figure 3: Generalisability use case. Left and centre, reconstructions for the EMPIAR-10278 dataset from the ground truth poses (grey) and the predicted poses (yellow) using a model trained on the EMPIAR-10280 data. Right, reconstruction for the EMPIAR-10280 dataset using a model trained on the EMPIAR-10280 data used as control (cyan). The model used for the yellow and cyan volumes is the same, but the particles fed to the model come from different datasets of the same protein in different conditions. Although of lower quality than the cyan volume, the yellow volume demonstrates similarities to the ground truth map (grey).
@@ 6 Acknowledgments
Appendix Appendix A Benchmark composition
@@
10399 Arabinofuranosyltransferase AftD from Mycobacteria C1 184 3.2 3.1 490616
10648 PKM2 in complex with Compound 5 D2 222 3.7 3.3 234956
simulated 10648 Same PKM2 dataset as in 10648, but with simulated images D2 222 3.5 3.4 138848
consensus 10648 Same PKM2 dataset as in 10648, but with consensus angles D2 222 3.8 3.4 234956
@@
Particle poses were estimated using the Relion version 4 auto-refine program[46, 24]. As starting model we used the map obtained with:
relion_reconstruct --pad 2.0 --ctf --i original_poses.star
--sym \$SYMMETRY --o reconstructd_map.mrc
The mask was created using:
relion_mask_create --i reconstructed_map.mrc --o mask.mrc --lowpass 15.0
-extend_inimask 3 --width_soft_edge 6 --ini_threshold \$THRESHOLD
with \$THRESHOLD manually selected for each entry.
The auto-refine command used was
mpirun -np 5 relion_refine_mpi --i original_poses.star --particle_diameter
\$DIAMETER --ctf --zero_mask --firstiter_cc --ini_high 40.0
--sym \$SYMMETRY --ref reconstructd_map.mrc --norm --scale
--solvent_mask mask.mrc --o outputdir/run --oversampling 1 --flatten_solvent
--solvent_correct_fsc --pad 2 --auto_local_healpix_order 4 --healpix_order 2
--offset_range 5.0 --offset_step 2.0 --auto_refine --split_random_halves
--low_resol_join_halves 40 --dont_combine_weights_via_disc
The simulated dataset was generated with the following command:
relion_project --i original_poses.star --ang original_poses.star
--ang_simulate original_poses.star --o simulated_dir/simulated
--simulate --adjust_simulation_SNR 2.0 --ctf
The consensus dataset was generated using the compare angles protocol from Scipion Xmipp [48, 21], incorporating both our original Relion refinement output and a refinement performed with cisTEM [16]. An angular distance threshold of 5° was employed.
@@
Notice that rotation transformations require adjustments in the ground truth labels.
@@
This section contains per-entry statistics for the Image2Sphere model predictions using the evaluation protocol proposed in the main text. The last two rows correspond to different versions of the 10648 entry and have not been included in the Table 1. In addition to angular error measurements, the other metrics compare the ground truth map ($\mathrm{GT}$) against the map reconstructed from the predicted poses ($\mathrm{V}$), namely the $\mathrm{PCC}(GT,V)$ and $\mathrm{FSCR_{t}}(GT,V)$, where $t$ denotes the threshold 0.5 or 0.143, where reported. GT is obtained by employing relion_reconstruct on the ground truth poses (that were estimated with relion_refine --auto_refine). The reconstructed map V is generated with relion_reconstruct from the predicted poses.
@@
10399 25.5 21.6 0.992 0.917 3.7 4.1 4.0 6.1 3.1 3.4 0.996
10648 13.3 10.6 0.982 0.934 3.8 4.1 4.3 6.5 3.4 3.6 0.994
simulated 10648 6.0 NA 0.996 0.935 4.5 4.6 4.6 4.8 3.5 4.6 0.998
consensus 10648 8.3 8.1 0.971 0.893 3.8 4.1 4.2 6.8 3.4 3.6 0.986
  • $\mathrm{MAnE}$: Mean Angular Error; $\mathrm{wMAnE}$: weighted Mean Angular Error; $\mathrm{PCC(V_{0},V_{1})}$: Reconstructed half-to-half Pearsons’s Correlation Coefficient; $\mathrm{PCC}(GT,V)$: Reconstructed to ground truth Pearson’s Correlation Coefficient; $\mathrm{FSCR_{0.143}(V_{0},V_{1})}$: Reconstructed half-to-half FSC resolution at threshold 0.143, and $\mathrm{FSCR_{0.5}(V_{0},V_{1})}$ at threshold 0.5; $\mathrm{FSCR_{0.143}(GT,V)}$: Reconstructed to ground truth resolution at threshold 0.143, and $\mathrm{FSCR_{0.5}}(GT,V)$ at threshold 0.5; $\mathrm{FSCR_{0.143}(GT_{0},GT_{1})}$: Ground truth half-to-half FSC resolution at threshold 0.143, and $\mathrm{FSCR_{0.5}(GT_{0},GT_{1})}$ at threshold 0.5. $\mathrm{PCC}(GT_{0},GT_{1})$: Ground truth half-to-half Pearsons’s Correlation Coefficient.
@@
This appendix shows the volumes reconstructed for some of the best performing examples of the Image2Sphere model on our benchmark. As is shown in all the cases, the quality of the central region of the protein is quite close to the one of the ground truth. However, the density for the regions that are at the edges of the macromolecule is much worse. This is in line to what could be expected if there were some degree of inaccuracy in the angular estimation, as the magnitude of the errors in the volume is proportional to both the angular error and the radius of the macromolecule.
Figure 1: Ground truth reconstruction for EMPIAR-10280 (grey) and reconstruction using the angles predicted with the Image2Sphere model (yellow).
Figure 2: Groundtruth reconstruction for EMPIAR-10374 (grey) and reconstruction using the angles predicted with the Image2Sphere model (yellow).
@@
In this section, we illustrate the usefulness of our approach by showing the effect of classical local refinement on the Image2Sphere results for the benchmark entry 10374. In this case, the Image2Sphere model predicted poses with a $\mathrm{wMAnE}$ of 24.8° that lead to a reconstructed map with $\mathrm{FSCR_{0.143}}(V_{0},V_{1})$ of 3.7 Å. When the predicted poses are used as priors for a local refinement in Relion with --sigma_angle 2.0, the refined map achieved a $\mathrm{PCC}(GT,V)$ of 0.997 compared to the original 0.969, showing that the refined map is much more similar to the ground truth map. Indeed, as it can be appreciated in Supplementary Figure 3, after the local refinement, not only the quality of the core of the protein is comparable to the quality of the ground truth, but also the quality of the distant parts of the maps is much better, almost as good as in the ground truth. Equally important, since we are limiting the angular search to the neighbourhood around the predicted poses ($\pm$ 6°), the number of image comparisons carried out by Relion is much smaller, resulting in a three-fold speed-up in computational time, even when including the time required for pose inference using the Image2Sphere model.
Figure 3: Ground truth reconstruction for EMPIAR-10374 (grey) and reconstruction using the angles predicted with the Image2Sphere model and locally refined using Relion with priors (yellow).
@@
Figure 5: Effect of particle misalignment in FSC resolution for entries 10648, 10786, and 11120. Left, FSC curves at different amounts of incorrectly aligned particles: 0% (blue), 10% (orange), 20% (green), and 30% (red) of the particles were assigned random angles. Right, FSC curves at different levels of induced angular inaccuracy. Each particle alignment was perturbed using uniform random noise of $\pm$0º (blue), 1º (orange), 3º (green), and 5º (red).